4-7x=1-6x

thanks
Creig

Answer by sdmmadam@yahoo.com(530)

4-7x=1-6x ----(1)
(-1+4)-7x+7x = (-1+1)-6x +7x [(adding (-1) from the left and 7x from the right to both the sides.]
3+0 = 0+x (using additive inverse law on the left and on the right)
3= x
That is x= 3
Verification: x = 3 in (1)
LHS = 4-7x = 4-7X3 = 4-21 = -17
RHS = 1-6x = 1-6X3 = 1-18 = -17 = LHS
Therefore our value x= 3 is right.